Problem 12-16 (LO. 2, 4) Seth, Pete, Cara, and Jen form Kingfisher Corporation with the following consideration: Consideratioif there is no gain or loss, enter $0 for the amount. As to these transactions, provide the following information for parts (e. Petes basis in the Kingfisher Corporation stock. $ 45,000 Feedback f. Kingfisher Corporations basis in the equipment. $h. Caras basis in the Kingfisher Corporation stock. $ 15,000 Feedback i. Kingfisher Corporations basis in the proprietary pI. During discussions relating to the formation of Kingfisher, Seth mentions that he may be interested in either (1) just sel
Answer #1

Answer to L since rest are correct:

Amount on sale = 90,000

Adjusted basis = 30,000

Ordinary gain = 90,000 - 30,000 = 60,000

Tax cost (15%) = 60,000 *15% = 9000

Present value of tax cost = (9000 * 0.7473 )+ (6000*35%) = 6726 + 2100 = 8826

Ans : $8,826